# A semi-analytical solution for the lubrication force between two spheres approaching in viscoelastic fluids described by the Oldroyd-B model under small Deborah numbers

### v1

@@ 5 Model predictions
While comparing with Zheng et al. (2023) provides insights into the behaviour of pressure and lubrication forces over time, it is typical to represent the lubrication force relative to the distance between the spheres’ contours. This can be achieved by evaluating (44) at various time intervals. In figure 8, results for a Newtonian fluid and for different viscoelastic fluids described by Oldroyd-B model are illustrated. Also, we compare our solutions to the predictions for a sphere moving towards a rigid plane in Ruangkriengsin et al. (2024) for $De=\{5\times 10^{-3},5\times 10^{-2}\}$. Ruangkriengsin’s predictions consider a Newtonian expression plus extra terms coming from the first-order and second-order corrections of the . The Newtonian fluid displays a linear trend, on logarithmic scales, along all values of $H_{0}$, whereas for the Oldroyd-B model a linear trend is exhibited only for values of $H_{0}$ between $\approx(0.1-0.6)$. In contrast, at $H_{0}=(0.6-1)$, the viscoelastic lubrication force begins at zero and increases until the Newtonian limit it reached. For shorter distances ($H_{0}\lesssim 0.1$), however, it surpasses the Newtonian lubrication force and continues increasing. Although comparing with Zheng et al. (2023) provides insight into the behaviour of pressure and lubrication forces over time, it is typical to represent the lubrication force relative to the distance between the spheres’ contours. This can be achieved by evaluating (44) at various time intervals. In figure 8, results for a Newtonian fluid and for different viscoelastic fluids described by Oldroyd-B model are illustrated. Also, we compare our solutions to the predictions for a sphere moving towards a rigid plane in Ruangkriengsin et al. (2024) for $De=\{5\times 10^{-3},5\times 10^{-2}\}$. Ruangkriengsin’s predictions consider a Newtonian expression plus extra terms resulting from first- and second-order corrections of the perturbation expansion. The Newtonian fluid displays a linear trend, on logarithmic scales, along all values of $H_{0}$, whereas for the Oldroyd-B model a linear trend is exhibit only for values of $H_{0}$ between $\approx(0.1-0.6)$. In contrast, at $H_{0}=(0.6-1)$, the viscoelastic lubrication force begins at zero and increases until the Newtonian limit it reached. For shorter distances ($H_{0}\lesssim 0.1$), however, it surpasses the Newtonian lubrication force and continues increasing. The enhancement of the viscoelastic lubrication force compared to the Newtonian lubrication force increases as distance reduces, as is detailed in table 1. Notably, even at the smallest De-number ($5\times 10^{-3}$), this increase can be as high as $70\%$, whilst at the higher De-number ($1\times 10^{-1}$), it can be up to two orders of magnitude greater than the Newtonian lubrication force. This considerable rise in lubrication force for viscoelastic fluids under transient dynamics, characterized by the Oldroyd-B model, could be crucial for modelling concentrated/dense systems with particles suspended in a viscoelastic medium with constant shear viscosity (Boger fluids). It is crucial to bear in mind that we have considered a constant De-number. However, as the spheres continue to close in, this gap changes and so does De, see (15), where De is expected to increase as the interparticle decreases, possibly further enhancing this effect.
@@
Figure 8: Dimensionless force against dimensionless interparticle distance for a Newtonian fluid and for different values of $De=\{5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},1\times 10^{-1}\}$ for $\beta=0$, and $\epsilon=2.5\times 10^{-3}$ and two estimates from Ruangkriengsin et al. (2024).
@@
Table 1: Percentage of the increase of the viscoelastic lubrication force against the Newtonian lubrication force for different De-numbers at different dimensionless interparticle distances.
@@ 5.1 Effect of the first normal stress-difference
To clarify the increase in the viscoelastic lubrication force using the Oldroyd-B model, we examined the topology of the velocity gradient of the interface separating both spheres. As outlined in Section 3, the viscoelastic response in the momentum balance is represented by the first normal stress-difference, such as: $\alpha(T)\left(\frac{\partial u}{\partial Z}\right)^{2}$, see (51), which denotes an extra resistance to shearing deformation of the fluid during the constant velocity approaching. Subsequently, this lubrication force enhancement may emerge by this extra term due to viscoelasticity, which scales with the square of the shear-rate. In addition, the major contribution to this resistance occurs near the centre of the particle gap. Figure 9 illustrates the velocity gradient magnitude in the interparticle separation between the spheres over time, where the blue colour represents low magnitude and red colour signifies high magnitude. When $H_{0}=0.49$, the velocity gradient across the space between particles remains minimal. At $H_{0}=0.10$, there is a slight increase in the velocity gradient magnitude on the surface of the spheres near the centre with magnitude around $\sim 1$, which leads to a minor rise in the lubrication force. As the distance decreases further, at $H_{0}=0.02$, the velocity gradient magnitude rises moderately to a pick value of $\sim 30$. Moreover, at distance $H_{0}=0.001$, the velocity gradient magnitude increase abruptly up to $\sim 3000$. Given that the effect of the viscoelasticity is proportional to the square of the velocity gradient magnitude, even with small values of De, the velocity gradient magnitude at $H_{0}=\{0.02,0.001\}$ is large enough to significantly enhance fluid resistance to be squeezed under constant velocity. This, in turn, increases the pressure gradient needed to sustain mass conservation. The result is a huge enhancement in the lubrication force near contact.
Figure 9: Dimensionless velocity gradient in the interparticle separation between the spheres for a viscoelastic fluid with $De=5\times 10^{-2}$, $\beta=0$, $a=1$, and $h_{0}=2.5\times 10^{-3}$.
@@ 5.2 Boger fluid problem
We consider the rheology data for the S1 fluid (a dissolution of polyacrylamide in corn syrup with composition: $0.05\%$ Separan AP 30, $2.4\%$ $H_{2}O$, and $97.55\%$ Corn Syrup 41N) in Phan-Thien et al. (1985), which may be described by Oldroyd-B model, in which the mechanics of a thin film being squeezed between two parallel plates is studied, in contrast to our current analysis, where two approaching spheres are considered. The flow properties of this Boger fluid are $\eta_{0}=55.5\;Pa\cdot s$, $\eta_{s}=15.8\;Pa\cdot s$ and $\lambda=5\;s$. In addition and for back reference, two Newtonian fluids are considered under $\eta_{0}=\{15.8,55.5\}\;Pa\cdot s$. The transient forces are calculated by evaluating (38) on the particle $S_{2}$ at different times under constant squeezing velocity. Moreover, we investigate the viscoelastic force when the spheres move away from each other. The approaching case starts from $h_{0}=2.5\times 10^{-4}$ to $h_{0}=2.5\times 10^{-7}$, meanwhile the moving away instance starts from $h_{0}=2.5\times 10^{-7}$ to $h_{0}=2.5\times 10^{-4}$, both cases kicking off from rest.
Figure 10 shows how the viscoelastic force behaves with varying De-numbers over time when utilizing the fluid S1 (Phan-Thien et al., 1985). As the spheres approach each other (illustrated in figure 10(a) as: solid lines for Newtonian fluids and filled symbols for Boger materials), the force increases as time progresses, whereas when the spheres separate (in figure 10(b) illustrated as: solid lines for Newtonian fluids and blank symbols for Boger fluids) the force decreases with time. For both Newtonian fluids, the lubrication force observed at the end of the approaching case match with the initial lubrication force noted during the separation scenario, and vice versa. Contrastingly, with the Boger fluids the lubrication force at the end of the approaching case is larger than the initial lubrication force at the beginning of the separation scenario. Consistently, the starting viscoelastic lubrication force during the separation starts at a lower level with respect to the end of the approaching case. During the sphere approach and separation modes, the viscoelastic lubrication force seems constrained by a lower limit corresponding to the Newtonian fluid of viscosity $\eta_{0}=\eta_{s}=15.8\;Pa\cdot s$, and an upper limit which corresponds to the Newtonian fluid of viscosity $\eta_{0}=\eta_{s}+\eta_{p}=55.5\;Pa\cdot s$. In the sphere approaching case, for $10.4\lesssim T<$ 1, the viscoelastic force exceeds the Newtonian reference, as predicted in Phan-Thien et al. (1985).
@@
In order to investigate how the evolution relates to distance, we apply the expression $H_{0}=1-T$ to calculate the dimensionless distance as time advances. figure 11 illustrates this situation for dimensionless distances from $H_{0}=1$ to $H_{0}=1\times 10^{-3}$, which corresponds to times from $T=0$ to $T=0.999$. For the sphere-approaching case, the lubrication force increases as the dimensionless distance decreases in time (in figure 11(a); here, the Boger fluid is illustrated with filled symbols, and time advances from right to left). In contrast, under the sphere-separating mode, the lubrication force decreases as distance increases with time (in figure 11(b); here, the Boger fluid is illustrated by blank symbols, and time advances from left to right). Notably and as a reference behaviour, Newtonian fluids follow the same trajectory under both approaching and separating modes (in figure 11(a) and 11(b), Newtonian cases illustrated by solid lines), displaying a symmetric response as a result of the Stokes conditions. Contrastingly, for the viscoelastic material in the approaching scenario, for distances between $H_{0}=1$ and $H_{0}=1\times 10^{-1}$, the viscoelastic force remains close to the upper Newtonian threshold. Conversely, when it drops below $H_{0}<1\times 10^{-1}$ at times $T>0.9$, there is a significant increase in the magnitude of the viscoelastic force reaching up to two orders-of-magnitude for $De=1.25\times 10^{-1}$. Under the sphere-separating instance, the viscoelastic case maintains its force near the lower Newtonian solution for $H_{0}<10^{-2}$. Beyond such landmark, a shift in the response occurs towards the upper Newtonian solution in the range $1\times 10^{-2}\leq H_{0}\leq 1\times 10^{-1}$. Hence, one may conclude that viscoelaticity may induce non-symmetric responses (see figure 12) and may be able to enhance the lubrication force during approaching case, even assuming very small De-numbers.
Figure 11: Dimensionless force on the particle against ddimensionless interparticle distance for the viscoelastic S1 fluid described in Phan-Thien et al. (1985) for different values of De-numbers and two Newtonian fluids. For viscoelastic Boger fluids, (a) the sphere-approaching mode is represented by filled symbols, whilst (b) the sphere-separating mode is identified through blank symbols. Newtonian fluids are illustrated by solid lines for both modes.
Figure 12: Scheme of the dynamic of the lubrication force for approaching and separating cases for two Newtonian lubrication forces (symmetric response drawn with continuous lines) and for a Boger fluid characterized by the Oldroyd-B model (non-symmetric response drawn with discontinuous lines).
@@ 6 Conclusions
In this work, the dynamics of the squeezing flow between two spheres embedded in a viscoelastic fluid, which may be described by an Oldroyd-B model, was examined. By focusing on the scenario of very narrow gaps between spheres, their shapes can be approximated as paraboloids. This simplification permits the use of lubrication theory to solve the mass and momentum balance equations employing an order-of-magnitude analysis. We consider weak viscoelasticity through small De-numbers. Differently to other approaches, we solved for the velocity profile analytically and employed numerical methods to calculate the pressure and normal force.  Once the squeeze flow is established, the momentum balance is composed by a viscous response coupled with a viscoelastic response leading by the first normal-stress differences (from the polymeric viscoelastic stress). Such momentum balance formulation reproduces a distinct velocity profile from the equivalently-viscous Newtonian profile, fact that evidences the viscoelastic contribution to the flow dynamics.  In the case of steady-state conditions, representing an immediate system response, viscoelasticity has a marginal effect in the lubrication force due to the small De-numbers imposed up to $\sim 2\%$. The lubrication force displays a linear increase with the De-number and $\beta$-fraction.  Conversely, for the transient solutions, changes of the gap lead to a time-dependent evolution of normal-stress difference that enhance resistance to flow by up to two orders of magnitude compared to the Newtonian response. This resistance is manifested through larger pressure gradients that work as the driven-force that ejects the fluid trapped between the spheres, even at the relatively small De-numbers imposed. Furthermore, viscoelasticity induces non-symmetric behaviour when spheres approach or separate. These phenomena arising from viscoelasticity may be utilised to modelling the rheological behaviour of dense suspension in viscoelastic Boger fluids, dominated by short-range interactions (Kumar et al., 2020; Ruiz-López et al., 2023; Rosales-Romero et al., 2024).
@@ Appendix A. Analytic solution for velocity profile
Here, the velocity profile derivation is given. The solution is calculated solving (51) using a reduction-order method, The Spanish State Research Agency through where $-\frac{dP^{\prime}}{dR}(T)=\psi$, $\alpha(T)=\alpha$, $\phi=\frac{du}{dZ}$, and $\frac{d\phi}{dZ}=\frac{d^{2}u}{dZ^{2}}$. Considering such definitions, (51) becomes:
@@
Applying the separation of variables, the solution of (58) is:
@@
By subtracting (60) from (61), an expression is obtained from which the first integration constant $C_{1}$ can be calculated, which is:
@@
From this results, the solution for (56) is:

### v2

@@ 5 Model predictions
Whilst comparison of our predictions against those by Zheng et al. (2023) provides insights into the behaviour of pressure and normal lubrication forces over time, it is typical to represent the lubrication force relative to the distance between the spherical contours. This can be achieved by evaluating (44) at various time intervals. In Figure 8, results are illustrated for a reference Newtonian fluid and for different viscoelastic fluids described by the Oldroyd-B model. The reference Newtonian fluid displays a linear declining trend with $H_{0}$-rise, whilst for the Oldroyd-B model, a linear trend is observed only for interpartle gaps in the range $H_{0}\approx(0.1-0.6)$. Notably, in the interval $H_{0}=(0.6,1]$, the viscoelastic lubrication force declines sharply to zero (see inset of Figure 8). In contrast, for shorter distances, i.e., $H_{0}\lesssim 0.1$, the viscoelastic lubrication force deviates from the Newtonian reference-line, increasing with $H_{0}$-decrease. This enhancement of the viscoelastic lubrication force with reducing interparticle gaps is detailed in Table 1. Notably, even at the smallest De-number examined, i.e., $De=5\times 10^{-3}$, this increase can be as high as $70\%$ over the Newtonian case, whilst for the largest De-number recorded, i.e., $De=1\times 10^{-1}$, this viscoelastic lubrication-force enhancement can amount over three orders-of-magnitude larger than the corresponding Newtonian instance, with a $2113\%$ excess. This rise in lubrication force for viscoelastic fluids under transient-confinement dynamics, evidenced here using the Oldroyd-B model, could be crucial for modelling dense systems with particles suspended in a viscoelastic medium with constant shear-viscosity Boger-fluid features. It is important to bear in mind that we have considered constant De numbers in our calculations. However, as spheres approach, the interparticle gap reduces, and, consequently, De increases (see (15)), possibly leading to further strengthening this viscoelastic enhancement on the normal lubrication force. In addition, we compare our solutions against the predictions for a sphere moving towards a rigid plane by Ruangkriengsin et al. (2024) for $De=\{5\times 10^{-3},5\times 10^{-2}\}$. Here, our predictions and those by Ruangkriengsin et al. (2024) concur in trends, only deviating at relatively reduced interparticle gaps and increased De numbers.
@@
Figure 8: Dimensionless viscoelastic lubrication force against interparticle distance under $De=\{5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},1\times 10^{-1}\}$, $\beta=0$, and $\epsilon=2.5\times 10^{-3}$ and two cases extracted from Ruangkriengsin et al. (2024) model.
@@
Table 1: Percent increase of the viscoelastic lubrication force over the Newtonian reference for a set of De-numbers and dimensionless interparticle distances.
@@ 5.1 Effect of the first normal stress-difference
To further describe the viscoelastic lubrication-force enhancement promoted by fluid confinement exposed in previous paragraphs, we examine the topology of the velocity gradient in the interstitial fluid separating both spheres. As outlined in Section 3, the viscoelastic component of stress in the momentum balance is represented by the first normal-stress difference, particularly contained in the term: $\alpha(T)\left(\frac{\partial u}{\partial Z}\right)^{2}$ (see (51) and the definition of the radial normal stress $S_{rr}$), which denotes an extra resistance to the shearing deformation of the fluid during the constant-velocity sphere-approaching modality. This lubrication-force enhancement may emerge through this extra term due to viscoelasticity, which scales with the square of the shear-rate. In addition, the major contribution to this viscoelastic resistance occurs near the centre of the particle gap. Figure 9 illustrates the velocity-gradient magnitude in the separation between the spheres over time; here, the blue colour-shading represents a relatively small shear-rate magnitude, whilst the red colour-shading depicts relatively large shear-rate values. Under the relatively large interparticle gaps of $H_{0}=0.49$, the velocity gradient across the space between particles remains minimal. At $H_{0}=0.10$, there is a slight increase in the velocity-gradient magnitude on the surface of the spheres near the centre, with magnitude around $\sim 1$ units, which leads to a minor rise in the lubrication force. As the interparticle gap decreases further in one order-of-magnitude up to $H_{0}=0.02$, the velocity-gradient magnitude rises moderately to a peak value of $\sim 30$ units. For the significantly-confined instance of $H_{0}=0.001$, the velocity-gradient magnitude increases abruptly up to $\sim 3000$ units. Given that the viscoelastic lubrication-force enhancement is proportional to the square of the velocity-gradient magnitude, even under small De-values, the velocity-gradient magnitude at $H_{0}=\{0.02,0.001\}$ is large enough to significantly enhance fluid resistance to be squeezed under a constant approaching velocity. This, in turn, and under such settings, increases the pressure gradient needed to comply with mass conservation, and results in a stark enhancement in the normal lubrication force near contact.
Figure 9: Dimensionless velocity gradient in the interparticle separation between the spheres for a viscoelastic fluid under $De=5\times 10^{-2}$, $\beta=0$, $a=1$, and $h_{0}=2.5\times 10^{-3}$.
@@ 5.2 Boger fluid problem
To further explore the predictive capabilities of our model, we consider the rheometrical data for the S1 fluid (a dissolution of polyacrylamide in corn syrup with composition: $0.05\%$ Separan AP 30, $2.4\%$ $H_{2}O$, and $97.55\%$ Corn Syrup 41N) in Phan-Thien et al. (1985). We characterise this fluid with the Oldroyd-B model, which yields the relevant material functions of $\eta_{0}=55.5\;Pa\cdot s$, $\eta_{s}=15.8\;Pa\cdot s$, and $\lambda=5\;s$. In addition and for back reference, two Newtonian fluids are considered under $\eta_{0}=\{15.8,55.5\}\;Pa\cdot s$. The transient forces are calculated by evaluating (38) on the particle $S_{2}$ at different times under a constant squeezing velocity. In addition, we investigate the viscoelastic force when the spheres move away from each other. The approaching case starts from $h_{0}=2.5\times 10^{-4}$ to $h_{0}=2.5\times 10^{-7}$, meanwhile the separating instance starts from $h_{0}=2.5\times 10^{-7}$ to $h_{0}=2.5\times 10^{-4}$, both cases kicking off from rest.
Figure 10 shows how the viscoelastic force behaves with varying De-numbers over time when utilising the fluid S1 (Phan-Thien et al., 1985). As the spheres approach each other (illustrated in Figure 10(a)), the lubrication force increases as time progresses, whilst when spheres separate (in Figure 10(b)) the lubrication force decreases with time. For both Newtonian fluids, the lubrication force observed at the end of the approaching case match with the initial lubrication force noted during the separation scenario, and vice versa. Contrastingly, with Boger fluids, the normal lubrication force at the end of the approaching case is larger than the initial lubrication force of the separation scenario. Consistently, the starting viscoelastic lubrication force during the separation starts at a lower level with respect to the end of the approaching case. During the sphere approach and separation modes, the viscoelastic lubrication force seems constrained by a lower limit corresponding to the Newtonian fluid of viscosity $\eta_{0}=\eta_{s}=15.8\;Pa\cdot s$, and an upper limit which corresponds to the Newtonian fluid of viscosity $\eta_{0}=\eta_{s}+\eta_{p}=55.5\;Pa\cdot s$. In the sphere approaching case, for $10.4\lesssim T<$ 1, the viscoelastic force exceeds the Newtonian reference, as reported in Phan-Thien et al. (1985).
@@
In order to investigate how the evolution relates to distance, we apply the expression $H_{0}=1-T$ to calculate the dimensionless distance as time advances. Figure 11 illustrates this situation for dimensionless distances from $H_{0}=1$ to $H_{0}=1\times 10^{-3}$, which corresponds to times from $T=0$ to $T=0.999$. For the sphere-approaching case, the lubrication force increases as the dimensionless distance decreases in time (in Figure 11(a); here, the Boger fluid is illustrated with filled symbols, and time advances from right to left). In contrast, under the sphere-separating mode, the lubrication force decreases as distance increases with time (in Figure 11(b); here, the Boger fluid is illustrated by hollow symbols, and time advances from left to right). Notably and as a reference behaviour, Newtonian fluids follow the same trajectory under both approaching and separating modes (in Figure 11(a) and 11(b), Newtonian cases are depicted by solid lines), displaying a symmetric response as a result of the Stokes conditions prevailing. Contrastingly, for the viscoelastic material in the approaching scenario, for distances between $H_{0}=1$ and $H_{0}=1\times 10^{-1}$, the viscoelastic force remains close to the upper Newtonian threshold. Conversely, when interparticle gaps drop below $H_{0}<1\times 10^{-1}$ at times $T>0.9$, there is a significant increase in the magnitude of the viscoelastic force rising up to two orders-of-magnitude for $De=1.25\times 10^{-1}$. Under the sphere-separating instance, the viscoelastic case maintains its force near the lower Newtonian solution for $H_{0}<10^{-2}$. Beyond such landmark, a shift in the response occurs towards the upper Newtonian solution in the range $1\times 10^{-2}\leq H_{0}\leq 1\times 10^{-1}$. Hence, one may conclude that viscoelaticity may induce asymmetric responses (see Figure 12) and may be able to enhance the lubrication force during approaching cases, even assuming small De-numbers.
Figure 11: Dimensionless force on the particle against interparticle distance for the viscoelastic S1 fluid described in Phan-Thien et al. (1985) for different values of De-numbers and two Newtonian fluids. For viscoelastic Boger fluids, (a) the sphere-approaching mode is represented by filled symbols, whilst (b) the sphere-separating mode is identified through blank symbols. Newtonian fluids are illustrated by solid lines for both modes.
Figure 12: Dynamics of the lubrication force for approaching and separating cases for two Newtonian lubrication forces (symmetric response, drawn with continuous lines) and for a Boger fluid characterised with the Oldroyd-B model (asymmetric response, drawn with discontinuous lines).
@@ 6 Conclusions
In this work, the dynamics of the squeezing flow between two spheres embedded in a viscoelastic fluid, described by an Oldroyd-B model, is examined. By focusing on the scenario of very narrow gaps between spheres, their shapes can be approximated as paraboloids. This simplification permits the use of lubrication theory to solve the mass and momentum balance equations employing an order-of-magnitude analysis. We consider weak viscoelasticity through small De-numbers. Differently to other approaches, we solved for the velocity profile analytically and employed numerical methods to calculate the pressure and normal forces.
Once the squeezing flow is established, the momentum balance is composed by a viscous response coupled with a viscoelastic response, led by the first normal-stress difference, originating from the polymeric viscoelastic stress. Such momentum balance formulation reproduces a distinct velocity profile from an equivalently-viscous Newtonian profile, fact that evidences the viscoelastic contribution to the flow dynamics.
We examine two flow regimes, i.e., approaching and departing sphere motion under steady and transient conditions. In the approaching case of steady-state conditions, representing an immediate system response, viscoelasticity has a marginal effect in the normal lubrication force due to the small De-numbers imposed. In this case, the normal lubrication force displays a linear increase with De-number and solvent fraction $\beta$.
Conversely, for transient solutions, changes in the interparticle gap lead to a time-dependent evolution of normal-stress differences that enhance resistance to flow by up to two orders-of-magnitude compared to the Newtonian response. This resistance is manifested through larger pressure gradients that work as the driving force that ejects the fluid trapped between the spheres, even at the relatively small De-numbers imposed. Furthermore, viscoelasticity induces asymmetric responses when spheres approach or separate. These phenomena arising from viscoelasticity may be utilised to modelling the rheological behaviour of dense suspension in viscoelastic Boger fluids, dominated by short-range interactions (Prasanna Kumar et al., 2020; Ruiz-López et al., 2023; Rosales-Romero et al., 2024).
@@ Appendix A. Analytic solution for velocity profile
Here, the velocity profile derivation is provided. The solution is calculated solving (51) using a reduction-order method, where $-\frac{dP^{\prime}}{dR}(T)=\psi$, $\alpha(T)=\alpha$, $\phi=\frac{du}{dZ}$, and $\frac{d\phi}{dZ}=\frac{d^{2}u}{dZ^{2}}$. Considering such definitions, (51) becomes:
@@
Under further integration, the solution of (58) is:
@@
By subtracting (60) from (61), an expression is obtained from which $C_{1}$ can be calculated, viz.:
@@
From these results, the solution for (56) is: